# Causal Lifting of Neural Representations: Zero-Shot Generalization for Causal Inferences

### v2

@@ Appendix A Proofs
  • Overlap Assumption, i.e., $10<\mathbb{E}[T|W=w]<$ 1 for all $w\in\mathcal{W}$,
  • Conditional Exchangeability (or Unconfoundness) Assumption, i.e., $\mathbb{P}(Y|do(T=t),W)=\mathbb{P}(Y|T=t,W)$ [^5] Re-framing it in do-calculus notation..
@@
Given a PPCI problem $\mathscr{P}$ with standard causal identification assumptions. If an outcome model $g:\mathcal{X}\rightarrow\mathcal{Y}$ is conditionally calibrated with respect to the treatment and a valid adjustment set, then the AIPW estimator  [Robins et al., 1994] over the prediction-powered sample with nuisance function estimators satisfying $||\hat{\mu}-\mu\|\cdot\|\hat{e}-e\|=o_{\mathbb{P}}(n^{-1/2})$, preserves the asymptotically valid confidence interval for the true ATE, i.e.,
$\sqrt{n}(\hat{\tau}_{g(X)}-\tau_{Y})\to\mathcal{N}(0,V),$
@@
Our goal is to prove that the following estimator for $\tau_{Y}$ is asymptotically normal:
@@
[…]
@@
Since we have from assumptions that $||\phi(\cdot;\hat{\mu}_{g(X)},\hat{e},g)-\phi(\cdot;\mu,e,g)||_{2}^{2}=o_{\mathbb{P}}(1)$, it holds that $T_{1}=o_{\mathbb{P}}(n^{-1/2}).$
@@
Given two similar PPCI problems $\mathscr{P}$ and $\mathscr{P}^{\prime}$ with standard causal identification assumptions. Let $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$ be an outcome model for $\mathscr{P}^{\prime}$ with $h$ Bayes-optimal, and assume that the representation transfers, i.e., $\mathbb{P}(\phi(X)|Y)=\mathbb{P}^{\prime}(\phi(X)|Y)$. Then the Causal Lifting constraint on $\mathscr{P}$ implies $g$ is valid on $\mathscr{P}$.
Intuition: The theorem shows that Causal Lifting, supported by other transferability assumptions, guarantees to transfer validity among PPCI problems, regardless of potential shifts in the joint distribution of $(Z,Y)$, e.g., different treatment effect. The core of the proof relies on establishing a hard representation transferability, i.e., $\mathbb{P}(\phi(X)|Y,Z)=\mathbb{P}^{\prime}(\phi(X)|Y,Z)$, given by the assumed representation transferability (on standalone $Y$) and the causal lifting constraint. This is done within the expectation of the conditional calibration, which we have already shown implies validity. Note that, indeed, the (marginal) representation transferability, conditioning over $Y$ alone, does not guarantee transferability conditioning over $Z$ too, potentially due by systematic biases within the encoder (e.g., fully retrieving the outcome signal for a certain subgroup and partially missing it for others).
@@ Appendix B ISTAnt
In this Section, we describe in detail our procedure to test zero-shot PPCI on ISTAnt dataset, by designing and recording a similar experiment for finetuning.
B.1 Similar experiment and data recording (ours)
We run an experiment very much alike the ISTAnt experiment with triplets of worker ants (one treated focal ant, and two nestmate ants), following the step-by-step design described their Appendix C [Cadei et al., 2024]. We recorded 5 batches of 9 simultaneously run replicates with similar background pen marking for the dish palettes positioning, producing 45 original videos, of which one had to be excluded for experimental problems, leaving 44 analyzable videos. Then, for each video, grooming events from the nestmate ants to the focal ant were annotated by a single domain expert, and we focus on the ‘or‘ events, i.e., “Is one of the nestmate ants grooming the focal one?”. We used a comparable experimental setup (i.e., camera set-up, random treatment assignment, etc.) except for the following, guaranteeing invariant annotation mechanism, i.e., similar experiment.
  • Treatments: Whereas ISTAnt used two micro-particle applications  [^6] By author correspondence., our experimental treatments also constitute micro-particle application in two different treatments ($n=15$ each), but also one treatment completely free of micro-particles (control, $n=14$), all applied to the focal ant. The three treatments of the ants are visually indistinguishable, independent of micro-particle application.
@@ B.1 Similar experiment and data recording (ours)
  • Longer Videos: Whereas ISTAnt annotated 10 min long videos, we here annotated 30 min long videos, even if the grooming activity of the ants decreases with time after exposure to treatment. Our videos were recorded at 30fps, totaling 158 400 annotated frames in the 44 videos.
@@
B.2 Zero-shot PPCI
@@ C.1 Data Generating Process
These data generating processes represent fictitious experiments where we aim to quantify the effect of the pen color on the number to draw (if asked to pick one), relying on a machine learning model for handwritten-digits classification. Particularly, the shift in the treatment effect between training (ATE$=1.5$) and target population (ATE$\in\{0,0.75,1.2\}$), may be interpreted as (i) testing a different, still homogeneous group of individuals, or (ii) varying the effect by changing the brand of the pens, unobserved variable (perfectly retrievable by the pen color on the training distribution), while keeping the same colors, reflecting the crucial challenges described in Section 2.1.
C.2 Analysis
@@
For each learning objective, i.e., DERM (ours), ERM, vREx and IRM, we trained a model on the training sample, imputed the outcome in each target sample and estimated the ATE via AIPW. Within the AIPW estimator we used the XGBoost regressor for the outcome regression and the XGBoost classifier to estimate the propensity score on observational studies (or vanilla sample mean on randomized controlled trials since constant). For both IRM and vREx we replicated the same hyper-parameter tuning of the invariant coefficient from our experiments on ISTAnt generalization, selecting in both case the invariant coefficient $\lambda=0.1$. We repeated each experiment 50 times, including resampling the data, and bootstrapped the confidence interval of the ATE estimates. We run all the analysis using $10\mathrm{GB}$ of RAM, $8$ CPU cores, and a single node GPU (NVIDIA GeForce RTX2080Ti). The main bottleneck of each experiment is re-generating a new version of CausalMNIST from MNIST dataset, and then training the model. The Prediction-Powered ATE estimation is significantly faster. We estimate a total of $12$ GPU hours to reproduce all the experiments described in this section.

### v3

@@ Appendix A Proofs
  • Overlap Assumption, i.e., $0<\mathbb{E}[T|W=w]<1$ for all $w\in\mathcal{W}$,
  • Conditional Exchangeability (or Unconfoundness) Assumption, i.e.,
    $\mathbb{P}(Y|do(T=t),W=w)=\mathbb{P}(Y|T=t,W=w)\qquad\forall w\in\mathcal{W},t\in\mathcal{T}.$
@@
Given a PPCI problem $\mathscr{P}$, with identifiable ATE (given the ground-truth annotations). If an outcome model $g:\mathcal{X}\rightarrow\mathcal{Y}$ is conditionally calibrated with respect to the treatment and a valid adjustment set, then the corresponding AIPW estimator over the prediction-powered sample with nuisance function estimators satisfying $||\hat{\mu}-\mu\|\cdot\|\hat{e}-e\|=o_{\mathbb{P}}(n^{-1/2})$, preserves doubly-robustness and valid asymptotic confidence intervals, i.e.,
$\sqrt{n}(\hat{\tau}_{g(X)}-\tau_{Y})\to\mathcal{N}(0,V),$ (14)
@@
Doubly-Robustness follows directly from the conditional calibration assumption and standard doubly-robust argument for AIPW  [Robins et al., 1994]. We then focus here on proving the asymptotically normality for $\tau_{Y}$ of the estimator:
@@
$\sqrt{n}\left(\frac{1}{n}\sum_{i=1}^{n}\phi(O_{i};\mu,e,g)-\tau_{Y}\right)\xrightarrow{d}\mathcal{N}(0,\underbrace{\mathbb{E}[\phi^{2}]}_{V}).$ (20)
@@
Since we have from assumptions that $\phi(\cdot;\hat{\mu}_{g(X)},\hat{e},g)-\phi(\cdot;\mu,e,g)||_{2}^{2}=o_{\mathbb{P}}(1)$, it holds that $T_{1}=o_{\mathbb{P}}(n^{-1/2}).$
@@
Given two similar PPCI problems $\mathscr{P}$ and $\mathscr{P}^{\prime}$, with identifiable ATEs given the ground-truth annotations. Let $g=h\circ\phi:\mathcal{X}\rightarrow\mathcal{Y}$ be an outcome model for $\mathscr{P}^{\prime}$ with $h$ Bayes-optimal, and assume that the representation transfers, i.e., $\mathbb{P}(\phi(X)|Y)=\mathbb{P}^{\prime}(\phi(X)|Y)$ for all $y\in\{\mathbb{P}(Y=y)>0\lor\mathbb{P}^{\prime}(Y=y)>0\}$. Then the Causal Lifting constraint on $\mathscr{P}$ implies $g$ is valid on $\mathscr{P}$.
Intuition: The theorem shows that Causal Lifting, supported by other transferability assumptions, guarantees to transfer validity among PPCI problems, regardless of potential shifts in the joint distribution of $(Z,Y)$, e.g., different treatment effect. The core of the proof relies on establishing a hard representation transferability, i.e., $\mathbb{P}(\phi(X)|Y=y,Z=z)=\mathbb{P}^{\prime}(\phi(X)|Y=y,Z=z)$, given by the assumed representation transferability (on standalone $Y$) and the causal lifting constraint. This is done within the expectation of the conditional calibration, which we have already shown implies validity. Note that, indeed, the (marginal) representation transferability, conditioning over $Y$ alone, does not guarantee transferability conditioning over $Z$ too, potentially due by systematic biases within the encoder (e.g., fully retrieving the outcome signal for a certain subgroup and partially missing it for others).
@@ Appendix B ISTAnt
In this Section, we describe in detail our procedure to test and achieve valid causal inferences on ISTAnt without human annotations but only complex measurements (video), relying on fine-tuning a foundational model on our similar annotated experiment.
B.1 Training annotated experiment and data recording (ours)
We run an experiment very much like the ISTAnt experiment with triplets of worker ants (one treated focal ant, and two nestmate ants), following the step-by-step design described in their Appendix C [Cadei et al., 2024]. We recorded 5 batches of 9 simultaneously run replicates with similar background pen marking for the dish palettes positioning, producing 45 original videos, of which one had to be excluded for experimental problems, leaving 44 analyzable videos. Then, for each video, grooming events from the nestmate ants to the focal ant were annotated by a single domain expert, and we focus on the ‘or‘ events, i.e., “Is one of the nestmate ants grooming the focal one?”, considering up to two possible behavior change per second (similarly for ISTAnt). We used a comparable experimental setup (i.e., camera set-up, random treatment assignment, etc.) except for the following, guaranteeing invariant annotation mechanism, i.e., similar experiment.
  • Treatments: Whereas ISTAnt used two micro-particle applications, our experimental treatments also constitute micro-particle application in two different treatments ($n=15$ each), but also one treatment completely free of micro-particles (control, $n=14$), all applied to the focal ant. The three treatments of the ants are visually indistinguishable, independent of micro-particle application.
@@ B.1 Training annotated experiment and data recording (ours)
  • Longer Videos: Whereas ISTAnt annotated 10-minute-long videos, we here annotated 30-minute-long videos, even if the ant grooming activity generally decreases with time from the first exposure to a new environment. Our videos were recorded at 30fps and analysed at 2fps, totaling 158 400 annotated frames in the 44 videos.
@@
B.2 Training and analyses details
@@ C.1 Data Generating Process
[…]
C.2 Additional Ablations: Causal Estimators
In addition to the main experiments on CausalMNIST we repeated all the experiments fine-tuning the predictive model via ERM and DERM, and varying the causal estimators. We particularly considered AIPW, X-Learner [Künzel et al., 2019], BART [Chipman et al., 2010], and Causal Forest [Wager and Athey, 2018]. Results are reported in Table 3. As expected, the same biases observed with AIPW are consistently confirmed by all the other estimators, confirming how biases in the prediction-powered samples invalidate any downstream analysis, regardless of the (causal) estimator [Cadei et al., 2024, Cadei and Internò, 2025].
Table 3: ATE bias with standard deviation over 50 repetitions on prediction-powered samples varying the causal estimators. Consistent biases are confirmed by different treatment effect estimators over the same prediction-powered sample.
Method Estimator In-Distribution OoD (linear effect) OoD (non-linear effect)
Soft shift Hard shift Soft shift Hard shift
ERM AIPW 0.00 $\pm$ 0.02 0.86 $\pm$ 0.14 1.05 $\pm$ 0.15 0.64 $\pm$ 0.18 0.85 $\pm$ 0.17
X-Learner 0.01 $\pm$ 0.02 0.82 $\pm$ 0.16 0.92 $\pm$ 0.15 0.68 $\pm$ 0.19 0.79 $\pm$ 0.18
BART 0.01 $\pm$ 0.02 0.81 $\pm$ 0.19 0.89 $\pm$ 0.13 0.59 $\pm$ 0.21 0.75 $\pm$ 0.16
Causal Forest 0.01 $\pm$ 0.02 0.82 $\pm$ 0.13 0.90 $\pm$ 0.11 0.64 $\pm$ 0.16 0.76 $\pm$ 0.12
DERM AIPW 0.10 $\pm$ 0.07 0.14 $\pm$ 0.14 0.75 $\pm$ 0.05 0.08 $\pm$ 0.27 0.45 $\pm$ 0.12
X-Learner 0.32 $\pm$ 0.27 0.00 $\pm$ 0.25 0.69 $\pm$ 0.12 0.05 $\pm$ 0.27 0.53 $\pm$ 0.14
BART 0.19 $\pm$ 0.14 0.07 $\pm$ 0.14 0.68 $\pm$ 0.15 0.06 $\pm$ 0.18 0.52 $\pm$ 0.18
Causal Forest 0.22 $\pm$ 0.19 0.11 $\pm$ 0.13 0.67 $\pm$ 0.09 0.03 $\pm$ 0.24 0.50 $\pm$ 0.15
C.3 Training and analyses details
@@
For each learning objective, i.e., DERM (ours), ERM, vREx, and IRM, we trained a model on the training sample, imputed the outcome in each target sample, and estimated the ATE via a causal estimator (AIPW for main experiments, AIPW, X-Learner, Causal Forest, and BART for the additional experiments). Within the AIPW and X-Learner estimator, we used the XGBoost regressor for the outcome regression and the XGBoost classifier to estimate the propensity score on observational studies (or vanilla sample mean on randomized controlled trials since constant). We rely on default parameters for both Causal Forest and BART (with 10 trees) as suggested in the original papers. For both IRM and vREx, we replicated the same hyper-parameter tuning of the invariant coefficient from our experiments on ISTAnt generalization, selecting in both cases the invariant coefficient $\lambda=0.1$. We repeated each experiment 50 times, including resampling the data, and bootstrapped the confidence interval of the ATE estimates. We run all the analysis using $10\mathrm{GB}$ of RAM, $8$ CPU cores, and a single node GPU (NVIDIA GeForce RTX2080Ti). The main bottleneck of each experiment is re-generating a new version of CausalMNIST from MNIST dataset, and then training the model. The Prediction-Powered ATE estimation is significantly faster. We estimate a total of $12$ GPU hours to reproduce all the experiments described in this section.